Find the axis of symmetry.
y = x^2 – x + 10
I am unsure of how to start this problem

Answer by stanbon(75887)

Find the axis of symmetry.
y = x^2 – x + 10
Rewrite as:
x^2-x+? = y-10+?
Complete the square, to get:
x^2-x+1/4 = y-10+1/4
Factor to get:
(x-(1/2))^2 = y-(39/4)
The axis of symmetry is x=1/2
Cheers,
Stan H.